X:Y = 5:3, Y:Z = 4:9
Find ratio X:Y:Z?

Answer by Edwin McCravy(20086)

X:Y = 5:3, Y:Z = 4:9
Find ratio X:Y:Z?

The idea is to multiply the terms of the ratios so as
to get the second term of the first ratio equal to the
first term of the second ratio.

The second term of the first ratio is 3 and the first
term of the second ratio is 4.

Now find the least common multiple of 3 and 4, which is 12.

So we multiply both terms of the first ratio by 4 and
both terms of the second ratio by 3.


X:Y = 5:3 = 5×4:3×4 = 20:12, 
Y:Z = 4:9 = 4×3:9×3 = 12:27

Therefore the ratio X:Y:Z = 20:12:27

Edwin